C3H8 + 5 O2 - - - - - - - - > 3 CO2 + 4 H2O
Calculate how many atoms of hydrogen in water are produced when 2.4 moles of propane are reacted
1
Expert's answer
2020-06-23T07:15:54-0400

"Reaction\newline C3H8\space+5O2\space \to\space 3CO2\space +4H2O.\newline Mole \space ratio \newline1:5\space\to3:4\space\space therefore,\newline No.\space moles\space of\space water(H2O)=2.4*4=9.6mol\newline In\space1mol\space of\space H2O\space there\space are\space 2\space hydrogen \space atoms\space so\space in \space 9.6mol\space we\space have\newline9.6mole*\dfrac{6.02*10^{23}molec}{1mol}*\dfrac{2atoms}{1molec}=1.15584*10^{25}"


"=""1.6*10^{25} \space atoms"